Find how many quarts of 6% buttermilk and 3% butterfat milk should be mixed yield 30 quarts of 4% butterfat milk.
The mixture should contain________ quarts of the 6% buttermilk
The mixture should contain _______ quarts of the 3% milk

Answer by jorel1380(3719)

x+y=30
.6x+.3y=30(.4)=12
6x+3y=120
3x+3y=90
3x=30
x=10
y=20
You need 10 qts. of 6% buttermilk and 20 qts. of 3% buttermilk..